For every 10% increase inprice we would predict a 25% decrease in quantity demand of goods. What type of elasticity is it (elastic or inelastic)

For every 15% decrease in price we would predict a 10% increase in quantity demand of goods. What type of elasticity is it (elastic or inelastic)?

Please show your calculation and explain your answer.
Answer #1

1. Price elasticity of Demand = % change in quantity demanded/ % change in price = 25/10 = 2.5

As the value of elasticity is greater than 1 therefore the demand is price elastic.

2. Price elasticity of demand = 10/15 = 0.67

As the value of elasticity is less than 1 therefore the demand is price inelastic.
Answer #2

To determine the type of elasticity in each scenario, we need to calculate the price elasticity of demand using the given information. Price elasticity of demand measures the responsiveness of quantity demanded to changes in price.

The formula for price elasticity of demand (PED) is:

PED = (% Change in Quantity Demanded) / (% Change in Price)

Let's calculate the price elasticity of demand in both scenarios:

Scenario 1: For every 10% increase in price, we predict a 25% decrease in quantity demand of goods.

% Change in Quantity Demanded = -25% % Change in Price = 10%

PED = (-25%) / (10%) = -2.5

Scenario 2: For every 15% decrease in price, we predict a 10% increase in quantity demand of goods.

% Change in Quantity Demanded = 10% % Change in Price = -15%

PED = (10%) / (-15%) ≈ -0.67

Now, let's interpret the results:

  1. Scenario 1: PED = -2.5

The magnitude of PED is greater than 1 (in absolute value), which indicates that the percentage change in quantity demanded is more than the percentage change in price. This means that demand is relatively responsive to price changes. The negative sign indicates an inverse relationship between price and quantity demanded, which is typical for most goods. This type of elasticity is known as elastic demand. In elastic demand, a change in price leads to a proportionally larger change in quantity demanded.

  1. Scenario 2: PED ≈ -0.67

The magnitude of PED is less than 1 (in absolute value), which indicates that the percentage change in quantity demanded is less than the percentage change in price. This means that demand is relatively unresponsive to price changes. The negative sign indicates an inverse relationship between price and quantity demanded. This type of elasticity is known as inelastic demand. In inelastic demand, a change in price leads to a proportionally smaller change in quantity demanded.

In summary:

  • Scenario 1 has elastic demand (PED = -2.5).

  • Scenario 2 has inelastic demand (PED ≈ -0.67).

Elasticity of demand plays a crucial role in pricing and revenue optimization strategies for businesses. Understanding the elasticity of demand helps companies make informed decisions on how to adjust prices and manage their sales and revenue targets.

answered by: Hydra Master